pooja195 (pooja195):

A test has twenty questions worth 100 points. The test consists of True/False questions worth 3 points each and multiple choice questions worth 11 points each. How many multiple choice questions are on the test?

Nnesha (nnesha):

test question are not allowed sorry :D

OpenStudy (loser66):

hahaha nope

pooja195 (pooja195):

-_- nnesha read the question xD

OpenStudy (loser66):

read the question green apple guy!!

pooja195 (pooja195):

*girl

OpenStudy (acxbox22):

set up da system of equations x+y=20 since the combination of both makes 20 questions and 3x+11y=100 since the tru/false is 3 points each and the multiple choice is 11 points each

pooja195 (pooja195):

thats a bit confusing :(

OpenStudy (alexandervonhumboldt2):

@Nnesha dont help if you dont want to

pooja195 (pooja195):

x + y = 20 3x + 11y = 100 is this right?

OpenStudy (acxbox22):

yes where x represents the number of t/f questions and y the number of multiple choice questions

pooja195 (pooja195):

Ah okay thanks :)

OpenStudy (acxbox22):

HOLD ON we arent done

OpenStudy (acxbox22):

now we do the solving

OpenStudy (acxbox22):

do u know how to solve system of equations?

pooja195 (pooja195):

3(x + y = 20) 3x + 11 y = 100 3x + 3y = 60 3x + 11y = 100 -8y = -40 8y = 40 y = 5

OpenStudy (acxbox22):

WOW very nice now plug in y in either equation and solve for x

pooja195 (pooja195):

There are 5 multiple choice questions. x + 5 = 20 x = 15 There are 15 T/F questions.

OpenStudy (acxbox22):

GOOD JOB! \(\Huge\color{green}{\ddot\smile}\)

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @AlexandervonHumboldt2 @Nnesha dont help if you dont want to \(\color{blue}{\text{End of Quote}}\) how do you know i don't wnat to help her :/

Nnesha (nnesha):

want*

pooja195 (pooja195):

Thanks so much i understand this :)